#5 hw Pensions   Meg's pension plan is an annuity with a guaranteed return of 7% per...

#5 hw

Pensions   Meg's pension plan is an annuity with a guaranteed return of 7% per year (compounded quarterly). She would like to retire with a pension of $30,000 per quarter for 25 years. If she works 41 years before retiring, how much money must she and her employer deposit each quarter? (Round your answer to the nearest cent.)

$
Answer #1

Given:
pension per quarter = $30,000
time=25 years
rate=7% compounded quarterly

Formulae to be used

FV = PV (1+i/n)(n*t)
   =30000*(1+0.07/4)(4*25)

   =170044.68
Hence, the future value she gets is 170044.68


Find how much shoud be deposited quarterly over 41 years to acquire future value.


Here, FV = 170044.67
i = 0.07
n=4
t = 41
Thus, the present amount is,

The present value can be calculated as:

FV = PV (1+i/n)(n*t)

==>PV = FV/ (1+i/n)(n*t)
           =170044.67/((1+0.07/4)(4*41))=$9883.68


Hence, the employee must deposit approx $9884 quarterly.

<<<<<<<<<<<<<<>>>>>>>>>>>>>>>>>>>>